Two complementary angles are 3r-4 degrees And r+30 degrees.Find the value of r.
Answer by math_tutor2020(3817)


Complementary angles form a 90 degree corner. I've highlighted the "co" to help remember the connection.

(3r-4)+(r+30) = 90
(3r+r)+(-4+30) = 90
4r+26 = 90
4r = 90-26
4r = 64
r = 64/4
r = 16